OpenStudy (anonymous):

obtain in polar form all values of z^pi z = 2+i

OpenStudy (anonymous):

\[z^{\pi}\]

OpenStudy (anonymous):

@dumbcow any idea ?

OpenStudy (dumbcow):

i think so... anyway you can rewrite z^pi in polar form as \[\large n^{\pi}(\frac{2}{n} +i \frac{1}{n})^{\pi} = n^{\pi}(\cos \theta +i \sin \theta)^{\pi}\] then solve for n and theta cos = 2/n sin = 1/n --> n = sqrt5 --> theta = .4636 radians so \[ \large (2+i)^{\pi} = \sqrt{5}^{\pi}(\cos (.4636 \pi) +i \sin(.4636 \pi))\]

OpenStudy (dumbcow):

not sure if angle has exact form....its arctan(1/2)

OpenStudy (anonymous):

okey i am checking

OpenStudy (dumbcow):

here is what wolfram says so it looks like i'm correct http://www.wolframalpha.com/input/?i=%282%2Bi%29^pi

OpenStudy (anonymous):

okey then :) thank u

OpenStudy (dumbcow):

no problem